OpenStudy (anonymous):

y=x-2 -x + y = -5 can anyone tell me how many solutions are in this problem

OpenStudy (nikato):

Change he second equation into the form y=mx+b

OpenStudy (anonymous):

can you just tell how many solutions are in this problem

OpenStudy (nikato):

No. I want u to find out by urself. I'm just gonna guide u

OpenStudy (anonymous):

thiis proble is so confusing

OpenStudy (nikato):

So change the second equation into y=mx+b

OpenStudy (anonymous):

how do i do that for this problem

OpenStudy (nikato):

So -x+y=-5 +x. +x ------------- Now what do u get after adding x to both sides?

OpenStudy (anonymous):

5x

OpenStudy (nikato):

No. It'll be y=x+5 And what can u tell me out the slope of this equation and the first one?

OpenStudy (anonymous):

thats it a no solution

OpenStudy (nikato):

Correct. Becuz both lines have the same slope and different y-intercepts, which means the lines are parallel and never meet

OpenStudy (anonymous):

and they just keep going

OpenStudy (nikato):

Correct and they will never meet

OpenStudy (nikato):

And no solution becuz u will never find a point that contains both lines